OpenStudy (anonymous):

2/y+3 + 3/y-4 = 5/y+6 Do you need to find the LCD?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so then I got 8/y+12 + -9/y+12 = 10/y+12 but I don;'t know where to go from there... if that is even correct?

OpenStudy (anonymous):

u have to multiply both the top and bottom, remember, what ever u do to the bottom u have to do the same to the top

OpenStudy (anonymous):

I don't know where to go from here though? 8/y+12 + -9/y+12 = 10/y+12

OpenStudy (shubhamsrg):

you probably have taken the LCD wrongly..do this,,instead of taking LCD,,multiply both sides by (y+3)(y-4)(y+6) this should help..

OpenStudy (anonymous):

I'm still lost by this? How would I do that?

OpenStudy (shubhamsrg):

like you did in the previous ques..whats wrong ?

OpenStudy (anonymous):

So 2+3(y+3)/y-4 = 5(y+3)/y+6

OpenStudy (anonymous):

does that look right at all?

OpenStudy (ajprincess):